OpenStudy (anonymous):

Solve for R: √2/2=24/R The Book says the answer is 24√2 my question is how they got this answer? i have no idea.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\sqrt{2}}{2} = \frac{24}{R}\] \[\Large R\sqrt{2}= 2*24\] \[\Large R\sqrt{2}= 48\] \[\Large R = \frac{48}{\sqrt{2}}\] \[\Large R = \frac{48\sqrt{2}}{\sqrt{2}*\sqrt{2}}\] \[\Large R = \frac{48\sqrt{2}}{2}\] \[\Large R = 24\sqrt{2}\]